OpenStudy (anonymous):

Which shows 63^2 - 37^2 being evaluated using the difference of squares method? 63^2 - 37^2 = (3969 + 1369)(3969 - 1369) = 13,878,800 63^2 - 37^2 = 3969 - 1369 = 2,600 63^2 - 37^2 = (63 - 37)^2 = 26^2 = 676 63^2 - 37^2 = (63 + 37)(63 - 37) = (100)(26) = 2,600 think its B

OpenStudy (anonymous):

The difference of squares method is using the fact that (a+b)(a-b) = a^2-b^2. In this case, you want to evaluate a^2-b^2, so you want to find out what a and b are, and plug those into the other side of the equation. a and b are clearly 63 and 37, so just plug them in: (63+37)(63-37) = (100)(26) = 2600 So, it's actually D.

OpenStudy (anonymous):

ohhh ! thanks, can you help me with another one please ?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Four polynomial functions are shown below: f(x) = x3 + 3x2 - 4x - 12 g(x) = x3 - 3x - 2 p(x) = x2 + 3x + 2 t(x) = x2 + x - 2 Which of the following functions has -3, -2, and 2 as its zeros? f(x) g(x) p(x) t(x)

OpenStudy (anonymous):

Actually, it's simple once you know how. If a polynomial has some number r as its root, then it has (x-r) as a factor. So, this polynomial has to have (x+3), (x+2), and (x-2) as factors. So, we just have to multiply these 3 factors. The last two give a difference of squares, (x^2-4), so it's (x+3)(x^2-4). You can just FOIL those to get x^3+3x^2-4x-12. And that's f(x), so there's your answer.

OpenStudy (anonymous):

oh i forgot about foil, thanks babe, 1 more pleaseee ;) ?

OpenStudy (anonymous):

Yeah, sure

OpenStudy (anonymous):

Which of the following best describes the graph of f(x) = x2 - 3x - 10? Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (anonymous):

Well, you've got two choices: find the minimum point, or find the roots. Finding the roots is just a matter of factoring. If you see that f(x) = (x-5)(x+2), then you have that the roots are 5 and -2, so it has intercepts at (5,0) and (-2,0). Alternatively, if you have ax^2+bx+c, the minimum is always at x = -b/2a. So, you would just do -(-3)/(2*1) = 1.5 for the x coordinate of the minimum. Either way, you'll end up with a.

OpenStudy (anonymous):

i have one more, it has to do with graphs, if you could help me that would be awesome

OpenStudy (anonymous):

Yeah, go ahead

OpenStudy (anonymous):

A polynomial function is shown below: f(x) = x3 - x2 - 9x + 9 Which graph best represents the function?

OpenStudy (anonymous):

Are there any other choices?

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

Good job, you're right. The easiest way to tell is to plug in 0, and find that y = 9. So, you need a graph with a y-intercept of 9, and the only one is C.

OpenStudy (the_fizicx99):

Lol Module 8 Part I

OpenStudy (anonymous):

thanks @srossd , finally a kind helping person,

OpenStudy (anonymous):

No problem, glad I could help.

OpenStudy (anonymous):

can i bother u with something else ?

OpenStudy (anonymous):

@srossd

OpenStudy (anonymous):

Yeah, go ahead

OpenStudy (anonymous):

Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (anonymous):

@srossd